### v2

# Faces of Fairness: Examining Bias in Facial Expression Recognition Datasets and Models

Abstract
Building AI systems, including Facial Expression Recognition (FER), involves two critical aspects: data and model design. Both components significantly influence bias and fairness in FER tasks. Issues related to bias and fairness in FER datasets and models remain underexplored. This study investigates bias sources in FER datasets and models. Four common FER datasets—AffectNet, ExpW, Fer2013, and RAF-DB—are analyzed. The findings demonstrate that AffectNet and ExpW exhibit high generalizability despite data imbalances. Additionally, this research evaluates the bias and fairness of eight deep models, including three generic convolutional neural network (CNN) models: MobileNet, ResNet, XceptionNet, as well as three generic transformer-based models: ViT, CLIP, and GPT-4o-mini, plus two FER-built state-of-the-art models: POSTER and CEPrompt. Experimental results reveal that among the 6 generic models (exclude POSTER and CEPrompt), GPT-4o-mini and ViT achieve the highest accuracy scores, while they also display the highest levels of bias. On the other hand, considering all the models demonstrate that the best accuracies are achieved for two FER-specialized models (POSTER and CEPrompt), while they could not cope with the bias problem in FER task. These findings underscore the urgent need for developing new methodologies to mitigate bias and ensure fairness in datasets and models, particularly in affective computing applications. See our implementation details at https://github.com/MMHosseini/bias_in_FER.
Index Terms:
Bias, Fairness, Facial Expression Recognition, Facial Affect Analysis, Bias in FER Datasets, Bias in FER Models.
1 Introduction
Facial expressions serve as a key non-verbal communication channel for humans [1], allowing them to express emotions and shape behavior, social interactions, and lifestyle [2]. Automated facial expression recognition (FER) has a borad range of applications, including mental health monitoring [3], human-computer interaction (HCI)[4], and surveillance[5]. Deep learning methods have recently achieved significant progress in FER, particularly through the use of convolutional neural networks (CNNs) [6], transformers [7], and large vision-language models  [8].
Despite advancements in deep learning, bias and fairness issues in FER systems remain unexplored. In a general overview, bias refers to prejudices or favoritism toward groups, while fairness ensures impartiality and avoids disparities [9, 10]. Bias can be analyzed in two dimensions: datasets and models [9, 11, 12]. Dataset bias often arises from demographic imbalances, such as disparities in age, gender, and race [9]. Model bias stems from architecture, training process, and unrepresentative datasets [13, 14, 15, 16]. Bias impacts model performance for underrepresented groups, reduces generalizability, and compromises fairness. Therefore, studying bias in datasets and models is crucial for fairness in FER.
There is a notable research gap in bias of automatic FER, particularly in in-the-wild datasets and novel deep models. In-the-wild datasets introduce biases from variations in illumination conditions, camera resolution, background noise, head pose, and gesture, as well as demographic attributes of age, gender, and race [17]. Many studies have focused on limited sources of bias in datasets, such as examining label distribution or one or two demographic attributes [18, 19, 20, 21], where they overlook the multifaceted nature of bias in FER. Additionally, recent advancements in deep architectures, such as vision transformers [22] and large vision-language models [23, 24], have significantly impacted this field. However, their implications for bias and fairness in FER tasks remain largely unexplored. These shortcomings prompted us to conduct an in-depth investigation into bias and fairness in facial expression recognition.
This research aims to address gaps by examining overlooked aspects of bias and fairness in FER. We provide a comprehensive analysis of bias and fairness in both datasets and deep models. Specifically, we argue that both datasets and models can introduce bias into FER systems. Fig. 1 shows the mathematical metrics that we utilized to evaluate various aspects of bias and fairness in the datasets and models.
To assess the impact of datasets, we analyze four real-world FER datasets: AffectNet [25], ExpW [5], RAF-DB [26], and Fer2013 [27]. The bias in these datasets is evaluated using seven well-known statistical metrics that measure diversity, uncertainty, and dependence in the data. We also investigate bias through analyses of single and joint probability distributions over the labels and demographic attributes of age, gender, and race. In addition, we examine inherent bias by conducting two model-based experiments: (1) identifying which dataset each sample originates from, and (2) evaluating how well models trained on three datasets generalize to a held-out dataset.
Figure 1: Our approach focuses on examining equity, including bias and fairness, in FER datasets and models. We leveraged various metrics to assess both datasets and models, where each metric quantifies specific aspects and enables numerical evaluation and comparison.
We study bias in models by evaluating eight popular deep models in FER. Specifically, we select three CNN models: MobileNetV1 [28], ResNet-50 [29], and XceptionNet [30]; three transformer-based models: ViT [22], CLIP [23], and GPT-4o-mini [24]; and two state-of-the-art FER-specialized models: POSTER [31] and CEPrompt [32]. We assess bias in these models using four key metrics that quantify how predictions or errors vary across demographic groups, capturing disparities in outcomes, opportunities, and relative treatment and evaluating overall fairness. By measuring multiple aspects of bias, these metrics provide a comprehensive evaluation across different attributes, allowing a fair and thorough comparison of the models’ bias scores. In summary, our main contributions on this research are:
  • Comprehensive bias assessment framework: a unified pipeline that analyzes both dataset and model bias in FER, with cross-dataset generalization tests and detailed demographic evaluations.
  • Advanced dataset-bias metrics: introduction of seven statistical and information-theoretic measures together with micro- and macro-level correlation analyses to reveal distributional and conditional dependencies.
  • Extensive model-bias benchmarking: fairness evaluation of eight representative CNN, Transformer, and FER-specialized models using four established fairness metrics, providing a reproducible benchmark for future studies.
Our implementation code and more detail about our experiments is reachable at GitHub.
In the remainder of this paper, Sec. 2 provides a review of the literature on bias and fairness in both datasets and algorithms. In continue, Sec. 3 outlines the metrics used for evaluating bias, while Sec. 4 discusses bias scores for the datasets and models. Finally, Sec. 5 concludes this bias study in FER.
2 Literature Review
Bias in machine learning can be classified in various ways. Pagano et al. [33] introduced three types of bias: pre-existing bias, technical bias, and emergent bias. Mehrabi et al. [9] provided a broader taxonomy including historical, representation, measurement, aggregation, and evaluation biases. Understanding the sources of bias is essential for effectively addressing its challenges. Bias can originate from both datasets and models. Factors such as data collection, cultural interpretation, annotation process, model architecture, objective functions, and evaluation methodologies are among the primary sources of bias. Developing fair and unbiased deep methods for FER potentially ensures equitable treatment across diverse demographic groups [34]. Thus, dealing with bias in FER has recently raised attention from researchers. In the following, we review recent studies on bias in FER datasets and models.
Figure 2: Bias can originate from datasets and models. In datasets, prominent sources of bias include issues in data collection, such as demographic disparities, variations in illumination and lighting conditions, gestures, head poses, and cultural differences in emotional interpretation. In models, key bias sources include architecture design, training parameters, overfitting to specific demographic groups, and the selection of evaluation metrics.
2.1 Bias in FER Datasets
Bias in datasets refers to systematic skews or unfairness in the data that can distort model training and predictions [35]. Several taxonomies have been proposed to categorize dataset bias. One common distinction is between representational bias and stereotypical bias[21, 36, 37]. The former targets demographic diversity, while the latter examines correlations between subgroups and labels. Another categorization separates intrinsic and extrinsic biases[38, 39, 40]: intrinsic bias arises from the data collection process and demographic or behavioral factors, whereas extrinsic bias stems from external conditions such as camera resolution, illumination, background noise, or annotator subjectivity. Additional dimensions have also been identified, with Cheong et al. [41] distinguishing among data collection bias, labeling bias, and contextual bias, and Jones et al. [42] categorizing bias into prevalence disparities, presentation disparities, and annotation disparities. Despite these diverse perspectives, there is strong consensus that demographic attributes, particularly age, gender, and race, remain the most influential sources of dataset bias and warrant the most attention.
Although establishing a taxonomy of dataset bias is important, understanding the underlying sources of bias provides guidance for mitigation. Based on our study, we categorize sources of bias into five main classes: 1) demographic disparities, referring to imbalances in age, gender, and racial groups that can affect model fairness across different populations; 2) environmental conditions, capturing inconsistencies in illumination, background scenes, and camera quality that may influence visual perception; 3) positional variations, including differences in body gestures, eye-gaze, head-pose, and clothing, which can alter the appearance of expressions; 4) expression diversity, reflecting imbalances between labels, and also easily recognizable and more subtle or challenging expressions that can bias model learning; and 5) annotation orientation, addressing cultural and subjective differences in labeling, which may introduce inconsistencies in ground truth data. A visual summary of these bias sources is provided in Fig. 2.
In the matter of in-the-wild FER datasets analysis, Dominguez et al. [21] found that while they are diverse in terms of demographic representation, they still exhibit weaknesses under stereotypical bias metrics. Li et al. [19] analyzed intrinsic bias across seven FER datasets, identifying cultural differences and data collection conditions as major contributors. Huber et al. [43] showed that synthetic facial datasets replicate the bias patterns present in the authentic datasets used for training. Cheong et al. [41] employed directed acyclic graphs to examine bias sources, highlighting gender and age as the most influential factors in emotion labeling. Collectively, these studies emphasize the critical role of dataset analysis in FER tasks, particularly the impact of demographic attributes on label distributions. Subsequent sections review common metrics for data analysis (Sec. 3.1) and present experimental results (Sec. 4.1).
2.2 Bias in FER Algorithms
Model bias is a systematic error in decision-making that produces unfair outcomes [12]. Another view describes it as biased results caused by inadequate model specifications [44]. Several taxonomies of bias in ML models have been proposed. Mehrabi et al. [9] identify algorithmic bias, evaluation bias, and optimization bias as key forms: algorithmic bias reflects the model’s assumptions, evaluation bias stems from benchmarks favoring certain groups, and optimization bias arises when objectives or loss functions benefit majority groups. Binns et al. [45] instead classify bias as direct, indirect, and structural. Direct bias occurs when algorithms deliberately treat groups unfairly to boost accuracy, indirect bias emerges from accidental errors when models fail to capture features, and structural bias results from broader organizational inequalities shaping algorithmic behavior.
Similar to dataset bias, we identify five main sources of bias in ML models: (1) algorithmic bias, which occurs when the model overfits to a dominant group in the training data; (2) model design bias, arising from choices in architecture, objective functions, loss functions, or training strategies; (3) feature extraction bias, when the model places undue emphasis on certain facial attributes; (4) generalization bias, where the model performs well only on a specific dataset but fails to transfer to others; and (5) evaluation bias, which stems from benchmarks or metrics that overlook fairness considerations (see Fig. 2).
Recent studies of bias in ML algorithms reveal that unfair outcomes in FER often persist even when data issues are addressed, suggesting that models themselves introduce bias. Dooley et al. [46] showed that bias can stem directly from network architecture, while Abdullah et al. [47] found that attention mechanisms allocate less focus to darker-skinned faces. Xu et al. [17] reported that data augmentation offers little improvement when the model is inherently biased, and Amigo et al. [34] demonstrated that models can reproduce bias across datasets despite debiasing strategies. Dominguez et al. [48] further argued that stereotypical biases are intrinsic to models, regardless of data balance. Together, these findings highlight that algorithmic bias is deeply embedded in model architectures and learning mechanisms, not just in the data.
2.3 Perspectives
Bias and fairness in facial expression recognition (FER) are essential for preventing discrimination, ensuring equity, improving performance across diverse groups, and maintaining ethical standards. Because real-world FER relies on in-the-wild datasets, it is crucial to examine these datasets to identify and address sources of bias. Although some studies have investigated them, a more systematic exploration with undiscovered metrics will open new insights into the inherent bias in the datasets. At the same time, rapid advancements in model architectures have significantly boosted FER accuracy [49, 50, 51]. Yet, despite these improvements, questions of fairness and bias in such models remain underexplored.
In this work, we conduct a comprehensive analysis of four widely used FER datasets (AffectNet, ExpW, Fer2013, and RAF-DB) using evaluation metrics. In parallel, we assess eight state-of-the-art models, including six general-purpose architectures (MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini) and two FER-specialized models (POSTER and CEPrompt), to investigate how bias evolves with model advancements. Taken together, this is the first systematic effort to provide an in-depth evaluation of both dataset- and model-level biases in FER.
3 Methodology
Bias and fairness are entangled and multifaceted concepts, with various definitions proposed in the literature (see [52]). Fairness is often described as the absence of favoritism or unjust privilege toward any group [9], while bias refers to prejudice or preferential treatment toward certain cases [21]. In essence, bias pertains to the data, whereas fairness focuses on the algorithms that process it. Numerous metrics have been introduced to measure both bias and fairness in datasets and algorithms (see [21, 53]). In the following sections, we delve into the selected metrics for evaluating datasets and algorithms, aiming to provide a comprehensive understanding of their impact on facial expression recognition.
TABLE I: Taxonomy of the equations used in this research.
Symbol Meaning Example
$A,B,C$ Symbol of each attribute AA: Age, $B$: Gender
$a,b,c$ Specific demographic group $a$: Man, $b$: Woman
$Y$ Expression set (labels) -
$y$ Specific expression value $y$: Happy
$H$ Entropy $H(Y)$: Entropy of labels
$n$ Number of labels (7 in our study) -
$k$ Number of demographic groups 4 age groups
METMET Models’ bias metrics -
ATTATT Attributes (age, gender, race) -
$p$ Probability -
3.1 Data Analysis
Bias in data manifest in ways, such as camera inconsistencies, background variations, body posture differences, and, most notably, demographic imbalances. Among these, demographic attributes, i.e., age, gender, and race, emerge as the most significant contributors to bias in FER datasets. As a result, these attributes, along with label distribution, warrant the most attention in this study to address bias and promote fairness.
To evaluate demographic bias in our datasets, we employ Wasserstein Distance (WD), Jensen–Shannon Divergence (JSD), Conditional-Entropy Bias Index (CEBI), Sparsity Index (SI), Normalized Shannon Entropy (NSE), Normalized Label Skewness (NLS), and Group Normalized Mutual Information (GNMI). Unlike the metrics in [21], these measures directly capture distributional differences and conditional dependencies that influence machine-learning models. WD and JSD quantify divergence in label distributions across demographic groups, revealing subtle shifts that broad diversity scores may miss. CEBI and GNMI measure information gain about labels from demographic attributes, making them sensitive to conditional or interaction-based bias. NSE and NLS offer normalized, interpretable indicators of entropy and skewness, enabling clear cross-dataset comparisons. We customized all the metrics in a way to work as a bias score in the range of $[0,1]$, where 0 means no bias and 1 shows the maximum bias.
Because our metrics draw on probability and information theory, they relate more directly to model training and predictive fairness than the ecological variety measures in [21]. This connection clarifies how demographic imbalance affects classifier performance and guides mitigation strategies. While classical diversity metrics help describe overall variety, the selected measures provide a complementary, machine-learning-focused view of dataset bias. Next, we present details for each metric. Table I presents the taxonomy of our equations.
Wasserstein Distance (WD), also known as Earth Mover’s Distance, is a metric sensitive to imbalance magnitude and class structure. For example, in a similar distribution of data for two attributes Male and Female, WD is equal to 0, while for two skewed and dissimilar distribution it gives a high value. This metric captures how much ”probability mass” should be shifted to reach to an equal data distribution:
[…]
Jensen–Shannon Divergence (JSD) quantifies the similarity between two probability distributions. While Wasserstein Distance quantifies bias by measuring the minimal effort required to transport one group’s distribution to another, Jensen–Shannon Divergence captures differences in their probabilistic information content without accounting for geometric distance. JSD is equal to 0 when two distributions are similar, and 1 when they are disjoint. In FER it determines how different the labels are distributed across demographic groups:
$\begin{aligned} JSD=\frac{2}{k(k-1)}\sum_{a,b\in A}\frac{1}{2n}\sum_{i=1}^{n}\left(H(a,m)+H(b,m)\right),\\ H(a,m)=p(y_{i}|A=a)\log\frac{p(y_{i}|A=a)}{m(y_{i})},\quad\quad\quad\\ m(y_{i})=\frac{p(y_{i}|A=a)+p(y_{i}|A=b)}{2}.\quad\quad\quad\quad\end{aligned}$
Conditional-Entropy Bias Index (CEBI) measures how an emotion class depends on a specific attribute in a group. For example, how the Male gender influences the Happy expression. A low CEBI means that the attribute strongly determines the label. To provide a consistent scale in which 0 indicates no bias and 1 indicates maximum bias, we normalized its equation and then reversed by subtracting it from 1:
$\begin{aligned} CEBI=\frac{1}{k}\sum_{a\in A}1-\frac{H(Y\mid a)}{H(Y)}.\end{aligned}$
Sparsity Index (SI) is a normalized version of Gini coefficient and quantifies the imbalance data distribution across the dataset in the range of [0, 1]. Unlike WD and JSD, which assess differences between group distributions, Sparsity Index measures how concentrated or uneven the label distribution is within each group, capturing intra-group imbalance rather than inter-group distance. For example, it evaluates how even the expressions are distributed on Male gender, independent of the Female gender. A higher SI reveals that only a few demographic groups dominate the samples, making it a clear indicator of representational bias:
$\begin{aligned} SI=\frac{1}{k}\sum_{a\in A}\frac{\left|\|\mathbf{p}\|_{2}^{2}-\tfrac{1}{n}\right|}{1-\tfrac{1}{n}},\quad\quad\quad\|\mathbf{p}\|_{2}^{2}=\sum_{y\in Y}p(y)^{2}.\end{aligned}$
Normalized Shannon Entropy (NSE) summarizes the entire demographic distribution into a single value without requiring pairwise group comparisons. NSE measures the overall label diversity within a group, providing a scale-invariant view of internal balance. For instance, it can evaluate how fairly the Happy expression is distributed across various racial groups; if it is biased toward an overrepresented group, NSE assigns a higher bias score. This metric simultaneously captures both the richness and evenness of group representation:
$\begin{aligned} NSE=\frac{1}{k}\sum_{a\in A}p(A=a)\left|1-\frac{H(Y\mid a)}{\log n}\right|.\end{aligned}$
Normalized Label Skewness (NLS) quantifies the asymmetry of class label distributions within each demographic group. For example, for the Male group, it measures how much each expression’s distribution deviates from the overall average. While Normalized Shannon Entropy (NSE) captures overall label diversity and balance, NLS highlights the directional imbalance of over- or underrepresented classes within a group. Large positive or negative values indicate strong bias, so NLS is normalized to the range [0, 1] for comparability:
$\begin{aligned} NLS=\frac{1}{k}\sum_{a\in A}\frac{|S|}{1+|S|},\quad\quad S=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{p(y_{i}|A=a)-\mu}{\sigma}\right]^{3}.\end{aligned}$
Group-Normalized Mutual Information (GNMI) measures non-linear statistical dependencies between labels and demographic groups. It quantifies how much information about a label can be inferred from group membership. For example, it evaluates the relationship between the Black racial group and the Happy expression, capturing how much knowing the demographic group reduces uncertainty about the label. While CEBI measures conditional label uncertainty given the group, GNMI normalizes this dependency for scale-independent comparison. High GNMI values indicate that the label is highly predictable from the demographic attribute:
$\begin{aligned} GNMI(Y,A)=\frac{I(Y;A)}{\sqrt{H(Y)H(A)}},\quad\quad\quad\\ I(Y;A)=\sum_{a\in A}\sum_{y\in Y}p(y,a)\log\frac{p(y,a)}{p(y)p(a)},\quad\end{aligned}$
Macro-Level Correlation Bias and Micro-Level Correlation Bias metrics examine the correlation between an attribute and a class label. Macro-Level Correlation Bias focuses on the overall relationship between an attribute and data distribution, such as the connection between ageage attribute and expressionexpression labels. On the other hand, Micro-Level Correlation Bias delves into analyzing the connection between a particular group and a class label. For example, the relationship between the group $age\mathord{=}[0\mathord{\sim}15]$ and $expression\mathord{=}Happy$.
To investigate Macro-Level Correlation Bias, we extracted the probability distributions of expressions and attributes in each dataset, denoted by $p(Y\mathord{=}y)$ and $p(A\mathord{=}a)$, respectively. To study Micro-Level Correlation Bias, given that each expression class is associated with three attributes—age, gender, and race—we explored the data distribution under single, double, and triplet conditions, denoted by $p(Y\mathord{=}y|A\mathord{=}a)$, $p(Y\mathord{=}y|A\mathord{=}a,B\mathord{=}b)$, and $p(Y\mathord{=}y|A\mathord{=}a,B\mathord{=}b,C\mathord{=}c)$, respectively.
We also evaluated non-demographic biases in each dataset. To do so, we designed an experiment where a model is trained to recognize the datasets. In this experiment, each image sample was labeled with its corresponding dataset name, for both the training and validation sets. Then, the model was trained to predict the dataset of each image. High accuracy in this experiment indicates higher non-demographic bias, such as those related to illumination, background noise, camera resolution, gesture, eye gaze, head pose, hair color, and any other hidden bias in the datasets.
In addition, we conducted another experiment to evaluate bias in datasets. In this leave-one-dataset-out experiment, we trained a deep model for facial expression recognition, using all except one dataset, which was left out for test. By assessing the model’s performance on the excluded dataset, we aimed to identify potential biases within the dataset. In this experiment, lower accuracy suggests higher bias in the datasets. To ensure a fair comparison, we trained the model on the training sets of all datasets except one, validated its accuracy using their respective validation sets, and tested the bias using the validation set of the excluded (left-out) dataset. If the results of the validation and test experiments show similar accuracies, the model is considered unbiased. Any discrepancy indicates that the dataset suffers from bias. To sum up, we utilized mathematical data analysis and designed two model-based experiments— dataset recognition model and leave-one-dataset-out FER model—to assess bias within the datasets.
3.2 Model Analysis
A variety of general metrics have been proposed to evaluate fairness in machine learning models [33, 9, 54].Verma et al. [52] introduced three categories of fairness parameters: statistical measures, similarity-based measures, and causal reasoning. Additionally, Mittal et al. [53] introduced metrics, such as parity-based metrics, score-based metrics, and facial-analysis-specific metrics.
While some researchers introduced their own metrics, most studies have concentrated on widely recognized metrics such as Equalized Odds [55], Equal Opportunity [55], Demographic Parity [56, 57], and Treatment Equality [58]. In this study, we focus on these four metrics to evaluate bias and fairness of models. To study bias in the models, we need to know the concept of protected and unprotected groups. Protected groups are privileged against discrimination, while unprotected groups may face exclusion or bias. In continue, we quickly review the four bias metrics we are going to utilize in this research.
Equalized Odds evaluates fairness by examining the positive predicted samples. Eq. 8 states that for an attribute $A$, a fair model will produce equal ratios of True Positives (TPs) and False Positives (FPs) for both the protected group ($A\mathord{=}a$) and unprotected group ($A\mathord{=}b$):
$\begin{aligned} p(\hat{Y}=1|A=a,Y=y)=p(\hat{Y}=1|A=b,Y=y).\end{aligned}$
Equal Opportunity focuses exclusively on TPs, where the probability of TPs ($Y\mathord{=}1$) for protected and an unprotected groups should be the same. The key difference between this metric and Equalized Odds is illustrated in Eq. 9. While Equalized Odds consider both TPs and FPs, Equal Opportunity concentrates on TPs:
$\begin{aligned} p(\hat{Y}=1|A=a,Y=1)=p(\hat{Y}=1|A=b,Y=1).\end{aligned}$
Demographic Parity examines the independence of predictions from attributes. In essence, it evaluates the distribution of positive predictions, irrespective of whether they are TP or FP. As Eq. 10 shows, this metric expects a uniform distribution of predictions across different groups:
$\begin{aligned} p(\hat{Y}=1|A=a)=p(\hat{Y}=1|A=b).\end{aligned}$
Treatment Equality examines the distribution of errors across different demographic groups (see Eq. 11). Specifically, it ensures that the ratio of False Negatives (FNs) to False Positives (FPs) remains consistent between protected and unprotected groups:
$\begin{aligned} \frac{p({\hat{Y}}=0|A=a,Y=1)}{p({\hat{Y}}=1|A=a,Y=0)}=\frac{p({\hat{Y}}=0|A=b,Y=1)}{p({\hat{Y}}=1|A=b,Y=0)}.\end{aligned}$
To obtain bias score of each model, for each demographic attribute, we applied all four bias metrics to each expression label. Next, we selected the maximum bias across all expressions as the bias score for that metric (see Eq.12). To assess the bias of each attribute, we averaged the bias scores across the different metrics using Eq.13. Finally, the overall bias of each model was calculated by averaging the bias score of each attribute (Eq. 14):
$\begin{aligned} Bias_{att}(met)=Max(Bias_{met,att}(y))~~~~\forall~y\in Y,\end{aligned}$
$\begin{aligned} Bias(att)=\frac{1}{4}\sum Bias_{att}(met)~~~~\forall~met\in MET,\end{aligned}$
$\begin{aligned} Bias=\frac{1}{3}\sum Bias(att)~~~~\forall~att\in ATT.\end{aligned}$
To sum up, we employed four well-established bias metrics to evaluate fairness of eight FER models over three demographic attributes, age, gender, and race. The detailed discussion of the models’ bias will follow in Section 4.
4 Experimental Results
This section explores bias across four well-known FER datasets: AffectNet [25], Fer2013 [27], RAF-DB [26], and ExpW [5]. Additionally, it examines fairness in eight generations of deep models, including three CNN models: MobileNet [28], ResNet [29], XceptionNet [30], as well as three transformer-based models: ViT [22], CLIP [23], and GPT-4o-mini [24], plus two state-of-the-art FER-specialized models: POSTER [31] and CEPrompt [32].
4.1 Bias in Datasets
Many studies suggest that the primary sources of bias in machine learning models lie in the data itself [10, 9, 11]. Consequently, to investigate bias in FER, it is essential to first examine the datasets. While numerous facial attributes can contribute to bias [59], and a variety of biases can be present within the datasets [41, 9], our focus is on the three most prominent facial attributes. In the following sections, we provide an overview of the datasets used in our study and detail the preprocessing steps we applied. Additionally, we explore the data distribution, attribute correlations, and model-based analysis.
4.1.1 Datasets
Although numerous FER datasets exist, only a few feature in-the-wild facial images. By considering factors such as popularity, number of expression labels, and annotation methods, we selected four datasets for our study: AffectNet [25], Fer2013 [27], RAF-DB [26], and ExpW [5]. All these datasets include at least six Ekman [60] basic expression labels, including Happy, Sad, Surprise, Fear, Disgust, Anger, plus Neutral, manually annotated by annotators.
TABLE II: Bias metrics for Age, Gender, Race, and their mean across datasets (in percent), include Wasserstein Distance (WD), Jensen-Shannon Divergence (JSD), Conditional-Entropy Bias Index (CEBI), Sparsity Index (SI), Normalized Shannon Entropy (NSE), Normalized Label Skewness (NLS), and Group Normalized Mutual Information (GNMI) metrics.
RAF-DB Fer2013 ExpW AffectNet
Age Gender Race Mean Age Gender Race Mean Age Gender Race Mean Age Gender Race Mean
WD 5.3 3.1 3.1 3.8 4.9 3.6 3.6 4.0 5.0 6.0 2.7 4.6 5.3 2.7 3.0 3.6
JSD 2.2 0.4 1.0 1.2 2.2 0.5 1.3 1.3 2.2 1.1 0.7 1.3 2.2 0.4 0.9 1.2
CEBI 82.2 63.7 90.5 78.8 83.1 64.8 89.4 79.1 82.0 61.5 91.0 78.2 81.6 62.5 90.3 78.2
SI 13.0 9.5 14.8 12.4 14.1 11.2 15.0 13.4 12.6 7.0 14.8 11.5 11.9 7.4 14.2 11.2
NSE 5.2 10.3 3.5 6.3 8.3 16.3 5.5 10.0 2.8 5.2 1.8 3.3 1.6 5.4 1.1 2.7
NLS 54.0 51.4 51.8 52.4 30.0 49.9 32.2 37.4 47.6 50.1 46.8 48.2 58.5 54.2 51.3 54.7
GNMI 61.9 29.3 76.0 55.8 54.7 26.1 70.2 50.4 64.4 31.7 83.7 59.9 67.1 33.3 84.3 61.6
Bias 30.1 27.9 29.6 30.5
AffectNet is the largest in-the-wild FER dataset, comprising nearly 1 million images, with half of them manually annotated. It includes eight expressions (seven basic expressions plus Contempt), continuous labels (valence-arousal), and additional metadata. Fer2013 contains approximately 36K images, divided into 29K training samples and 3.5K validation samples. However, its small image size ($\times40$ pixels) and grayscale color representation pose limitations. RAF-DB provides both single-label and compound-label subsets, with around 30K facial images and facial landmark annotations. The compound-labeling approach in RAF-DB results in 19 expressions, including the seven basic expressions and 12 compound expressions such as Happily-Surprise and Sadly-Angry. Finally, ExpW is a large-scale FER dataset, featuring over 90K manually annotated facial images. We utilize these datasets in our study to train and evaluate our models.
[…]
4.1.2 Data Preprocessing
To have a unique form of data, in the first step, we aligned images by ensuring the same cropping area, image size, and color domain across all the datasets. We extended the original crop region by 25 to 35 percent in each dimension for all images, then used the DeepFace model [61, 62] to process the extended crop and align the facial images. Next, we resized all the cropped faces to $224\times 224$. Since Fer2013 is grayscale, we replicated the grayscale layer three times to match the RGB format, ensuring uniformity across the data. Additionally, ExpW lacks a validation set, therefore, we randomly selected 10 percent of its images to serve as the validation set.
4.1.3 Metric-Based Dataset Analysis
[…]
In the initial experiment, we assessed dataset bias using the metrics described in Sec. 3.1. By applying Eqs. 1 to 7, we derived the bias scores summarized in Table II. This table presents the scores for each metric and demographic attribute, where the final bias score of each dataset is the average of different scores.
The bias analysis of the four datasets reveals notable demographic disparities, with overall bias scores ranging from 27.9% (Fer2013) to 30.5% (AffectNet). None of the datasets are fully balanced across Age, Gender, and Race, suggesting that models trained on these datasets are likely to inherit these imbalances. Among the attributes, Race consistently exhibits the highest bias values across most metrics, indicating that racial representation is more uneven than Age and Gender. Gender shows comparatively lower bias, although metrics like CEBI and NLS indicate persistent disparities, while Age-related bias is moderate but consistent. These findings highlight that demographic representation is far from uniform, and differences in attribute distribution could affect fairness in model predictions.
Examining individual metrics provides deeper insight into the nature of these biases. Metrics such as CEBI and GNMI report high values across all datasets, indicating strong conditional dependencies and unequal mutual information between demographic attributes and labels. SI values reveal the uneven distribution of samples across different attribute-label combinations, while NSE and NLS reflect deviations from uniform label distributions, with Fer2013 showing higher unevenness in Age and Gender, and AffectNet showing pronounced skew in Age. WD and JSD are relatively lower, suggesting that while overall distributional differences exist, they are less pronounced than the conditional dependencies captured by CEBI and GNMI. Overall, these results underscore the importance of considering both dataset composition and metric-specific insights when addressing bias in facial expression recognition.
[…]
4.1.4 Attributes Distribution
[…]
As shown in Fig. 3-c, the number of Man samples is approximately 50% more than Woman, though this gender bias is less pronounced compared to other attributes. Fig. 3-d, on the other hand, highlights a notable racial imbalance, with White samples constituting over 50% of the total across all datasets. The distribution of Asian, Black, Latinx, and Middle Eastern samples is more balanced, while Indian samples remain consistently underrepresented. In conclusion, an examination of the data distribution across four essential attributes—age, gender, race, and expression—uncovers notable bias in the datasets, where addressing them is vital to ensure fairness in FER models.
4.1.5 Attributes Correlation
Building on Micro-Level Correlation Bias, we studied the joint probability distribution of data. Examining correlations between joint attributes provides a meaningful approach for data analysis. We kept expression as the primary attribute and analyzed its correlation with other facial attributes: age, gender, and race.
[…]
Figure 5: The 4D correlation matrix between different attributes of all the datasets is presented, with expressions represented in the rows and the columns divided first by age groups, followed by gender, and finally by race groups. This heat map highlights an uneven data distribution, where a great portion of data are underrepresented. This diagram reveals limited and imbalance diversity in the datasets.
[…]
Figure 6: The confusion matrix for FER datasets (in percent) reveals that the high diagonal values for Fer2013, ExpW, and AffectNet point to inherent biases within these datasets. Notably, RAF-DB stands out as the only dataset that shows a significant correlation with the others.
4.1.6 Dataset Generality
Based on the data distribution and correlation analysis, we examined demographic biases within the datasets. While demographic disparity is a primary source of bias, other factors such as illumination, head pose, image quality, background settings, lighting effects, and facial accessories also contribute to bias in dataset [39, 40, 64]. These sources of bias vary significantly across different datasets.
To investigate general bias in FER datasets, we conducted an experiment where a model was trained to identify the dataset to which any given image belongs. Each image was labeled with its respective dataset name, and an XceptionNet model [30] was trained to predict the dataset label. The results, as shown in Fig. 6, demonstrated biases inherent to FER datasets. While some correlation exists between RAF-DB and two other datsets (ExpW and AffectNet), Fer2013, ExpW, and AffectNet remain uncorrelated. This experiment underscores that differences in data collection methodologies—including image acquisition, processing, cropping, and storage—significantly affect FER datasets. Such biases, stemming from dataset-specific protocols, remain a relatively underexplored drawback in FER research.
TABLE III: Bias analysis across different datasets (in percent). When one dataset is excluded, a model is trained on the remaining datasets. Model is evaluated on the validation set of the included datasets, and then is tested over the excluded dataset. More difference between validation and test scores highlights more inherent biases in trained datasets.
RAF-DB Fer2013 ExpW AffectNet
Validation 59.0 54.1 59.0 57.0
Test 60.1 71.2 44.0 45.9
An additional experiment was conducted to further investigate the generality and biases inherent in the datasets. In this experiment, one dataset was excluded, and a facial expression recognition model (XceptionNet) was trained using the remaining datasets. To ensure fairness, the model was first evaluated on the unseen validation sets of the datasets included in training. Subsequently, the generalizability of the excluded dataset was assessed by testing the model on its images. Since the excluded dataset was not part of the training process, the results reflect the dataset’s bias. Table III compares the validation accuracies achieved on the included datasets with the test accuracy for the excluded dataset. Interestingly, test accuracies for RAF-DB and Fer2013 were higher than their validation accuracies. This indicates that two other datasets, ExpW and AffectNet used for training, have provided good generalability. An opposite trend was observed for ExpW and AffectNet which shows that when we used RAF-DB and Fer2013 as train sets the model could not capture challenhes in ExpW and AffectNet datasets. For instance, when AffectNet was excluded, validation accuracy across the validation sets of RAF-DB, Fer2013, and ExpW was 57.0%, while test accuracy on the excluded AffectNet dataset dropped to 45.9%. This experiment gives us clues for more inherent bias over RAF-DB and Fer2013 datasets, and lower bias over ExpW and AffectNet.
4.1.7 Conclusion
The analysis of the four FER datasets reveals that all exhibit notable demographic and expression biases. None of the datasets are fully balanced across Age, Gender, Race, or expression categories, and these imbalances may affect model performance and fairness. While the degree and nature of bias vary slightly between datasets, the overall patterns are consistent, highlighting the importance of considering dataset composition when developing and evaluating facial expression recognition models.
4.2 Fairness of Models
After analyzing bias in the FER datasets, we further investigate bias and fairness within the models. As mentioned, four bias metrics are used to evaluate eight deep models with different architectures. In the following sections, we describe our training methodology and discuss the accuracy and bias of each model.
Figure 7: Model accuracy for each expression (in percent). In this experiment, models were trained on individual datasets and evaluated on their respective evaluation sets. $\mu$ and $\sigma$ denote the average accuracy and its standard deviation for each dataset, while MEAN and STD represent the overall average accuracy and standard deviation across each model. RAF-DB emerged as the least challenging dataset, whereas ExpW posed the most difficulty. Among the models, POSTER reached the highest accuracy, contrasting with CLIP which showed the lowest accuracy.
4.2.1 Training and Implementation Details
We selected eight deep learning architectures, including six basic CNN- and transformer-based models: MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini, plus two SOTA FER-specialized models: POSTER and CEPrompt. MobileNet was used as a baseline CNN model due to its streamlined feed-forward structure without residual connections, making it computationally efficient for mobile and embedded applications. ResNet and XceptionNet, incorporating residual connections, were chosen for their ability to mitigate the vanishing gradient problem and improve model accuracy. Among the transformer-based models, ViT represented a state-of-the-art approach for computer vision tasks, leveraging self-attention mechanisms instead of traditional convolutional operations. CLIP, a large vision-language model designed for multi-modal learning, enabled joint processing of images and text. GPT-4o-mini, a lightweight variant of GPT-4, was included for its capability to efficiently handle both NLP and vision tasks. To better understand bias in FER, we examined the POSTER model, which integrates self-attention with posterior distribution modeling to capture subtle emotion cues, and the CEPrompt model, which applies context-aware prompts to guide vision transformers for more robust, context-sensitive facial expression recognition.
The experiment was conducted using Python 3.8.10, TensorFlow 2.9.2 [65], and the Hugging Face [66] libraries. Four Nvidia 1080 GPUs with 8 GB and 12 GB of memory were utilized. As a preprocessing step, image dimensions were fixed at $224\times 224$. We conducted experiments on each dataset and model separately and also merged the training and validation sets of all datasets for subsequent experiments. Notably, after merging the datasets, we obtained 404,755 training samples and 22,858 validation samples.
In the first experiment, we evaluated each model’s accuracy on every dataset separately. In the second, we merged all training sets, trained each model, and then measured accuracy on the combined validation set. Finally, the third experiment focused on assessing each model’s bias utilizing the bias metrics. For this, we fully trained four models—MobileNet, ResNet, XceptionNet, and ViT—and fine-tuned CLIP, POSTER, and CEPrompt. Full training was conducted for 30 epochs, while fine-tuning was limited to 5 epochs. We used the Adam optimizer with a learning rate of $10^{-3}$ for MobileNet, ResNet, XceptionNet, CLIP, and CEPrompt, and $10^{-5}$ for ViT and POSTER. Categorical cross-entropy served as the loss function for all training procedures.
Since GPT-4o-mini is not open source, we evaluated it without any fine-tuning. We prompted all the facial images of the validation sets into this model and asked it to assign one of the seven expressions used in this study to that image. The prompt we used was: ”What is the expression of this person among happy, sad, surprise, fear, disgust, anger, and neutral, in only one word? Try to find one of these expressions, but if your prediction is out of these expressions, select the closest expression to it (among the 7 aforementioned expressions). If you cannot determine any expression, randomly select one of the mentioned expressions. If the predicted expression is another word in the family of the aforementioned expressions (happy, sad, surprise, fear, disgust, anger, neutral) change it to the corresponding word. For example, fearful is equal to fear, disgusted is equal to disgust, angry is equal to anger, surprised is equal to surprise, scared is equal to fear, and so on. If you randomly select an expression, do not need to explain, only generate the expression name.
4.2.2 Facial Expression Recognition
Figure 7 shows the accuracy of each model across the datasets, categorized by expression. The results highlight a noticeable bias toward Neutral and Happy expressions, where they achieved significantly higher accuracy scores compared to the other expressions, in all the models. When comparing the expression distribution in Figure 4 with the accuracy results in Figure 7, it becomes evident that data imbalance contributes to this bias. Nonetheless, the models displayed an ability to partially compensate for this imbalance. For example, although the ratio of Disgust to Happy data samples was highly skewed ($\frac{2.2}{41.8}=0.05$), the accuracy gap between these two expressions was considerably narrower ($\frac{29.1}{85.9}=0.33$), suggesting that the models made efforts to mitigate this bias during training.
Figure 8: Confusion matix of each model across all datasets (in percent). In this experiment, all the datasets were used for both of the training and validation stages. No model was able to achieve high scores for the two challenging expressions, Fear and Disgust. A comparison of the results reveals that GPT-4o-mini and ViT models demonstrated the most consistent accuracy scores, while CLIP struggled to achieve satisfactory performance.
The average accuracy for each dataset, represented by $\mu$ in Fig. 7, demonstrates a correlation between dataset richness and complexity. Richer datasets tend to be more challenging, leading to lower average accuracies. The last two columns of this figure, labeled Mean and STD, summarize the mean accuracy and standard deviation for each model. Through this figure, POSTER and CEPrompt (two FER-specialized models) achieved the highest accuracy, which looks normal. Among the rest of 6 basic models, GPT-4o-mini achieved the highest accuracy, with minimum standard deviation. While this suggests POSTER, CEPrompt and GPT-4o-mini may be the fairest models, subsequent experiments reveal they exhibit significant bias.
In contrast, the CLIP model showed the worst performance, with an average accuracy of 49.7% and a high standard deviation exceeding 25%. These results indicate that the CLIP model is not well-optimized for FER tasks in its current form. Training the entire model, rather than just fine-tuning its decision-making head, could potentially improve its performance. Lastly, among the fully trained models (MobileNet, ResNet, XceptionNet, and ViT), ViT achieved the highest accuracy score, while ResNet exhibited the lowest standard deviation. Based solely on high accuracy and low standard deviation metrics, the models can be ranked from least to most accurate models as follows: POSTER, CEPrompt, GPT-4o-mini, ViT, ResNet, MobileNet, XceptionNet, and CLIP. We will discuss this ranking in Sec. 4.2.4 when we extract the metric-based bias scores.
Another experiment for FER models involved training each model using the training samples from all datasets and evaluating their performance on all validation sets. Fig. 8 displays the confusion matrix for these general models across all validation sets. At first glance, it is clear that Fear and Disgust are the most challenging expressions for all models, with none achieving more than 50% accuracy in predicting them.
TABLE IV: General accuracy of each model (in percent): four models—MobileNet, ResNet, XceptionNet, and ViT—were fully trained; three models—CLIP, POSTER, and CEPrompt—were fine-tuned; and GPT-4o-mini was evaluated solely on the validation sets without fine-tuning.
Neu Hap Sad Sur Fea Dis Ang Mean STD
MobileNet 66.6 86.8 53.2 60.1 45.6 31.0 57.5 57.3 17.4
ResNet 73.8 84.1 48.0 60.2 38.2 24.4 54.0 54.7 20.4
XceptionNet 77.3 82.3 51.5 61.3 44.8 24.7 57.7 57.1 19.6
ViT 75.3 86.7 56.5 64.0 42.8 25.1 64.3 59.2 20.4
CLIP 79.7 85.1 44.1 45.1 25.8 9.4 43.8 47.5 25.8
GPT-4o-mini 72.6 79.5 56.8 62.9 27.6 34.9 62.1 56.6 17.5
POSTER 79.9 89.1 62.9 64.4 48.2 29.1 63.5 62.4 18.2
CEPrompt 79.3 88.8 61.7 68.4 42.4 29.2 62.7 61.7 18.9
TABLE V: Bias analysis using Equalized Odds, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
Neu Hap Sad Sur Fea Dis Ang Max Mean STD
Age 4.9 2.1 1.1 0.3 1.0 2.0 1.1 4.9 1.7 1.3
Gender 7.7 12.7 2.2 2.3 0.1 5.9 2.8 12.7 4.8 3.9
MobileNet Race 11.6 9.0 7.1 5.9 2.2 7.7 4.7 11.6 6.8 2.8
Age 6.7 1.8 0.7 0.4 1.7 0.6 2.1 6.7 1.9 2.0
Gender 6.9 11.6 0.8 2.3 0.5 2.8 4.4 11.6 4.1 3.6
ResNet Race 10.1 6.3 7.1 4.7 1.4 5.0 3.4 10.1 5.4 2.5
Age 5.1 2.0 0.9 0.5 1.9 1.0 1.1 5.1 1.7 1.4
Gender 7.2 12.2 2.1 2.5 0.9 3.2 2.9 12.2 4.4 3.6
XceptionNet Race 10.2 7.9 6.3 5.7 1.4 5.2 2.8 10.2 5.6 2.7
Age 4.7 2.0 0.5 0.7 0.8 1.4 2.8 4.7 1.8 1.3
Gender 9.5 13.4 0.5 2.9 0.4 2.8 4.8 13.4 4.8 4.4
ViT Race 13.1 9.6 9.1 6.0 1.5 4.3 5.2 13.1 6.9 3.5
Age 7.3 2.8 0.1 0.9 2.7 0.1 2.6 7.3 2.3 2.3
Gender 8.4 11.0 1.3 2.2 1.2 0.1 6.7 11.0 4.4 3.9
CLIP Race 12.5 7.4 7.4 4.9 4.7 1.1 5.6 12.5 6.2 3.2
Age 3.3 1.3 2.2 0.5 1.2 0.6 4.1 4.1 1.8 1.2
Gender 6.1 9.4 1.7 3.1 1.2 1.4 9.7 9.7 4.6 3.4
GPT-4o-mini Race 10.3 8.1 5.4 6.2 2.2 3.3 9.7 10.3 6.4 2.8
Age 6.4 1.1 0.4 0.5 2.1 0.4 2.8 6.4 1.9 2.0
Gender 5.9 10.6 1.2 2.0 1.5 0.4 6.6 10.6 4.0 3.4
POSTER Race 10.8 6.9 6.9 4.8 5.6 2.1 6.3 10.8 6.2 2.4
Age 5.1 1.0 0.1 0.6 2.0 0.3 3.0 5.1 1.7 1.6
Gender 7.3 10.2 1.7 1.8 1.6 0.2 6.9 10.2 4.2 3.5
CEPrompt Race 12.4 6.3 6.1 5.4 4.4 2.1 6.4 12.4 6.1 2.9
Deeper analysis of Fig. 8 reveals significant confusion among certain expressions. For instance, pairs like Sad-Neutral, Disgust-Neutral, and Disgust-Anger pose considerable challenges, with erroneous predictions in the range of $18.8\%-35.3\%$, while zero values would be expected. The expressions accuracy below 50% are highlighted by red color in this figure. A comparison between POSTER and CEPrompt models and the rest of the models illustrates that they could not tackle the problem of Fear and Disgust expressions in FER task. On the other hand, while Fig. 7 showed, excluding two FER-specialized models, GPT-4o-mini as the top performer across various datasets, the general model results, summarized in Table IV, indicate that the ViT model achieved the highest average accuracy in this experiment. Additionally, MobileNet had the lowest standard deviation, indicating more consistency in its predictions.
TABLE VI: Bias analysis using Equal Opportunity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
Neu Hap Sad Sur Fea Dis Ang Max Mean STD
Age 8.3 2.0 4.0 5.2 4.6 0.4 1.5 8.3 3.7 2.4
Gender 1.8 11.8 4.4 4.4 1.1 7.7 4.6 11.8 5.1 3.3
MobileNet Race 3.2 16.5 17.9 20.5 6.8 14.4 4.8 20.5 12.0 6.4
Age 9.5 3.0 1.8 4.9 3.8 2.3 0.7 9.5 3.7 2.6
Gender 1.2 9.5 0.4 4.5 3.8 2.2 9.0 9.5 4.3 3.3
ResNet Race 5.6 14.9 21.3 21.2 8.8 15.1 2.9 21.3 12.8 6.7
Age 8.5 4.3 0.6 6.8 6.0 1.3 3.1 8.5 4.3 2.6
Gender 1.2 10.9 0.8 4.9 2.6 9.0 8.5 10.9 5.4 3.7
XceptionNet Race 6.1 15.7 16.1 19.2 10.4 11.0 5.7 19.2 12.0 4.7
Age 5.8 1.6 4.3 3.7 4.4 0.1 1.4 5.8 3.0 1.8
Gender 4.3 11.9 7.1 8.3 3.6 5.5 11.0 11.9 7.3 2.9
ViT Race 4.6 17.1 24.9 25.2 11.1 9.6 10.4 25.2 14.7 7.3
Age 7.3 1.7 3.1 9.4 6.6 4.4 0.2 9.4 4.6 3.0
Gender 2.8 9.0 2.8 7.9 7.4 7.4 15.4 15.4 7.5 3.9
CLIP Race 6.5 12.0 21.4 15.4 27.3 7.6 13.0 27.3 14.7 6.8
Age 4.5 0.0 6.5 5.8 0.6 9.6 2.7 9.6 4.2 3.1
Gender 1.0 6.9 3.0 10.7 13.6 14.9 17.9 17.9 9.7 5.8
GPT-4o-mini Race 4.9 15.3 11.2 28.9 14.9 11.5 10.5 28.9 13.8 6.9
Age 5.6 0.5 0.4 5.0 1.9 1.3 0.2 5.6 2.1 2.0
Gender 1.6 7.8 1.4 8.9 10.2 6.0 11.6 11.6 6.7 3.7
POSTER Race 6.0 8.2 12.1 24.3 25.0 19.2 8.0 25.0 14.6 7.4
Age 2.8 0.3 2.8 7.9 0.4 5.6 0.4 7.9 2.8 2.7
Gender 5.3 6.8 1.7 5.9 8.5 7.2 11.6 11.6 6.7 2.8
CEPrompt Race 8.8 8.5 11.0 24.2 20.0 9.5 6.6 24.2 12.6 6.1
So far, we have assessed the models’ accuracy to evaluate their robustness against bias stemming from imbalanced data distribution. Our findings indicate that, overall, POSTER and CEPrompt are the most accurate models, while CLIP performed poorly. Although these results suggest that POSTER, CEPrompt, GPT-4o-mini, and ViT might be the least biased models, subsequent experiments reveal different outcomes. In the next step, we will apply various bias metrics to further analyze the models’ biases across different attributes.
4.2.3 Metric-Based Model Analysis
Using the expression results from the previous experiments and the age, gender, and race attributes extracted with DeepFace [63], we evaluated the bias of all eight models with the metrics introduced in Sec. 2.2.
The first metric, Equalized Odds, examines the fairness of True Positive (TP) and False Positive (FP) rates across different groups. Table V highlights that Neutral and Happy were the most biased expressions. Almost all maximum biases for demographic attributes originate from these two expressions. For the Age attribute, Surprise emerged as the least biased expression, whereas for Gender and Race attributes, Fear showed the lowest bias. The Max column of this table reveals that the Age attribute consistently exhibited the lowest bias across all models. The Mean and STD columns further emphasize that, on average, Race was the most biased attribute across all models, while Age demonstrated the least bias. When comparing the Mean and STD values among the models, the Equalized Odds metric indicates that the ViT model was the most biased, while ResNet and XceptionNet demonstrated the lowest bias scores.
TABLE VII: Bias analysis using Demographic Parity, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
Neu Hap Sad Sur Fea Dis Ang Max Mean STD
Age 5.0 0.9 0.9 0.3 1.0 3.1 1.1 5.0 1.7 1.5
Gender 8.4 15.6 0.7 2.3 0.1 7.4 2.9 15.6 5.3 5.1
MobileNet Race 9.1 8.0 6.9 5.4 1.9 8.1 2.9 9.1 6.0 2.5
Age 6.5 0.8 0.8 0.1 1.7 1.0 2.2 6.5 1.8 1.9
Gender 7.3 13.9 1.0 2.3 0.1 3.1 4.5 13.9 4.6 4.3
ResNet Race 7.2 7.6 7.0 4.6 0.8 4.4 3.4 7.6 5.0 2.2
Age 5.1 0.3 1.3 0.4 2.0 1.3 1.2 5.1 1.6 1.5
Gender 8.6 14.9 2.6 2.4 0.5 3.4 3.2 14.9 5.0 4.6
XceptionNet Race 10.0 9.5 6.8 5.4 1.6 4.8 2.6 10.0 5.8 2.9
Age 4.6 0.1 0.4 0.5 0.9 1.6 2.8 4.6 1.5 1.5
Gender 11.3 16.1 0.1 2.8 0.3 3.1 5.0 16.1 5.5 5.5
ViT Race 13.5 9.9 9.0 6.4 1.2 3.7 5.3 13.5 7.0 3.8
Age 6.5 1.8 0.2 1.1 2.7 0.1 2.9 6.5 2.1 2.0
Gender 10.6 13.9 0.0 2.5 0.9 0.1 6.8 13.9 4.9 5.1
CLIP Race 11.9 5.9 7.6 4.2 4.6 1.2 4.5 11.9 5.7 3.0
Age 3.6 0.6 1.6 0.6 1.3 0.4 4.4 4.4 1.7 1.4
Gender 8.4 12.8 0.1 3.3 1.2 0.6 9.5 12.8 5.1 4.6
GPT-4o-mini Race 5.5 9.5 5.5 6.8 2.3 2.9 9.8 9.8 6.0 2.7
Age 6.0 0.7 0.6 0.7 2.0 0.3 3.1 6.0 1.9 1.9
Gender 8.9 14.1 1.8 2.7 1.2 0.4 6.9 14.1 5.1 4.6
POSTER Race 12.3 7.2 5.5 5.4 5.3 2.0 4.5 12.3 6.0 2.9
Age 4.9 0.7 0.2 0.9 1.8 0.2 3.2 4.9 1.7 1.6
Gender 9.8 13.7 0.3 2.4 1.3 0.2 7.0 13.7 4.9 4.9
CEPrompt Race 13.2 7.0 5.8 5.8 4.4 1.6 5.7 13.2 6.2 3.2
Another metric we introduced in Sec.3.2 is Equal Opportunity, which focuses on the ratio of TP predictions across different groups. Information about this metric is provided in Table VI. Unlike Equalized Odds, this table shows that the Happy expression was not the most biased in terms of Equal Opportunity. For the Age attribute, Neutral ranked highest in bias, while for Gender and Race attributes, Happy and Surprise expressions exhibited the highest bias scores, respectively. Examining the Max, Mean, and STD columns of this table reveals that Race is the most bias-affected attribute across all models. Based on this metric, GPT-4o-mini exhibited the highest bias, while ResNet and MobileNet achieved the lowest bias scores.
Demographic Parity analyzes the likelihood of any group receiving a particular outcome. The data in Table VII shows that, compared to other bias metrics, Demographic Parity results in lower bias scores. For the Age attribute, the highest bias was observed in the Neutral expression, while Happy, Sad, and Surprise had minimal effects. In contrast, Happy was the most biased expression for the Gender attribute, while Fear and Sad showed the least bias. For the Race attribute, Neutral and Happy were the primary sources of bias, while Fear exhibited the lowest bias across all models. Although the Max column highlights higher biases for the Gender attribute, the Mean column underscores Race as the most biased attribute. As with the other bias metrics, Age remained the least biased attribute. The Mean and STD columns illustrate that for the Age attribute, the CLIP model was the most biased model, while for Gender and Race, the ViT model was the most biased model. Overall, considering the Mean bias scores, the ViT model emerged as the most biased, while ResNet demonstrated the least bias.
TABLE VIII: Bias analysis using the Treatment Equality, in the range of [0, 100], across models and three demographic attributes: age, gender, and race. Bias is calculated for each expression, followed by the calculation of maximum value, mean, and standard deviation for each attribute.
Neu Hap Sad Sur Fea Dis Ang Max Mean STD
Age 1.1 1.3 7.9 6.6 5.2 3.5 0.3 7.9 3.7 2.7
Gender 1.9 26.4 12.4 14.3 1.6 8.7 6.5 26.4 10.2 7.9
MobileNet Race 28.4 46.8 38.8 19.9 21.7 11.0 54.3 54.3 31.5 14.4
Age 8.4 3.3 11.0 11.3 14.1 13.6 5.1 14.1 9.5 3.8
Gender 11.4 10.4 6.5 12.8 9.3 1.2 16.0 16.0 9.6 4.3
ResNet Race 49.7 28.0 23.8 14.2 44.0 32.8 25.4 49.7 31.1 11.3
Age 1.5 11.4 0.6 1.7 14.5 17.3 1.5 17.3 6.9 6.6
Gender 3.0 22.5 10.3 12.3 1.5 5.4 0.2 22.5 7.8 7.2
XceptionNet Race 35.7 48.2 14.5 24.9 33.9 29.8 19.1 48.2 29.4 10.4
Age 0.7 3.9 1.0 1.6 3.4 24.2 11.5 24.2 6.6 7.9
Gender 7.7 34.8 2.8 4.9 1.0 16.0 16.2 34.8 11.9 10.8
ViT Race 29.5 61.7 44.8 22.8 19.5 31.4 43.5 61.7 36.1 13.6
Age 6.5 8.3 6.3 2.0 23.5 0.0 1.8 23.5 6.9 7.3
Gender 1.0 15.2 4.9 1.1 1.7 3.2 33.6 33.6 8.6 11.1
CLIP Race 8.8 23.2 29.0 29.1 26.3 4.4 9.4 29.1 18.5 9.8
Age 6.2 6.3 15.8 1.9 4.2 14.1 5.1 15.8 7.6 4.8
Gender 5.3 8.2 9.0 7.7 1.6 43.5 45.3 45.3 17.2 17.3
GPT-4o-mini Race 48.4 61.0 13.5 46.6 26.6 27.0 39.8 61.0 37.5 14.9
Age 7.4 5.9 12.2 1.0 13.3 6.1 5.2 13.3 7.3 3.9
Gender 1.4 31.9 20.5 6.5 6.4 0.5 14.3 31.9 11.6 10.5
POSTER Race 10.7 37.1 24.3 30.7 37.5 12.2 39.4 39.4 27.4 11.1
Age 1.6 6.4 6.1 4.9 12.7 4.6 3.4 12.7 5.6 3.2
Gender 8.7 23.1 6.3 17.8 8.3 8.6 21.7 23.1 13.5 6.5
CEPrompt Race 14.4 37.3 20.9 46.0 37.3 16.5 34.8 46.0 29.6 11.2
The final bias measure we examined is Treatment Equality, which assesses the distribution of misclassified samples across different groups. Table VIII demonstrates that Treatment Equality resulted in higher bias scores. A detailed expression-level analysis shows that Sad, Fear, and Disgust were the main sources of bias for the Age attribute. Happy exhibited the highest bias for the Gender and Race attributes.Among the attributes, Race had the highest bias scores, while Age exhibited the lowest. This trend was consistent with our findings in the Equalized Odds and Equal Opportunity metrics. Comparing the Max column in Table VIII with the Max columns in Tables VVI, and VII shows that Treatment Equality exhibited higher bias scores across all attributes. In a model-wise comparison of Treatment Equality, CLIP showed the lowest bias score, while the highest bias was observed in ViT.
TABLE IX: Bias score for all models using the measures Equalized Odds (Eq-Od), Equal Opportunity (Eq-Op), Demographic Parity (De-Pa), and Treatment Equality (Tr-Eq) across age, race, and gender attributes. The mean represents the average of each row, and the bias score is the final mean score of the attributes. Data are in the range of [0, 100].
Eq-Od Eq-Op De-Pa Tr-Eq Mean Bias Score
Age 4.9 8.3 5.0 7.9 6.5
Gender 12.7 11.8 15.6 26.4 16.6
MobileNet Race 11.6 20.5 9.1 54.3 23.8 15.6
Age 6.7 9.5 6.5 14.1 9.2
Gender 11.6 9.5 13.9 16.0 12.7
ResNet Race 10.1 21.3 7.6 49.7 22.1 14.6
Age 5.1 8.5 5.1 17.3 9.0
Gender 12.2 10.9 14.9 22.5 15.1
XceptionNet Race 10.2 19.2 10.0 48.2 21.9 15.3
Age 4.7 5.8 4.6 24.2 9.8
Gender 13.4 11.9 16.1 34.8 19.0
ViT Race 13.1 25.2 13.5 61.7 28.3 19.0
Age 7.3 9.4 6.5 23.5 11.6
Gender 11.0 15.4 13.9 33.6 18.4
CLIP Race 12.5 27.3 11.9 29.1 20.2 16.7
Age 4.1 9.6 4.4 15.8 8.4
Gender 9.7 17.9 12.8 45.3 21.4
GPT-4o-mini Race 10.3 28.9 9.8 61.0 27.5 19.1
Age 6.4 5.6 6.0 13.3 7.8
Gender 10.6 11.6 14.1 31.9 17.0
POSTER Race 10.8 25.0 12.3 39.4 21.8 15.5
Age 5.1 7.9 4.9 12.7 7.6
Gender 10.2 11.6 13.7 23.1 14.6
CEPrompt Race 12.4 24.2 13.2 46.0 23.9 15.4
To sum up, we analyzed bias of the models for each attribute and expression, followed by a model-based bias comparison. Based on our experiments, ViT showed the highest bias in the Equalized Odds, Demographic Parity, and Treatment Equality metrics, while GPT-4o-mini exhibited significant bias in Equal Opportunity. On the other hand, ResNet was the most robust model against bias in three out of four bias metrics.
Investigating bias scores for FER-specialized models, POSTER [31] and CEPrompt [32], reveals that despite achieving high accuracies in FER task, they suffer bias problem. Base on the ranking of models in each bias metric they mainly are located between the third to sixth place out of 8 models, which is not considerable. A peer-to-peer comparison between POSTER and CEPrompt showed that in three out of four metrics CEPrompt showed lower bias than POSTER.
Table IX presents the final bias score of each model across different attributes and bias metrics. Based on the Eq. 12, we used the maximum bias score of each model as its bias score. In the next step, using the Eq. 13, we calculated the average bias score for each model across all attributes. Finally, through the Eq.  14, the overall bias score for each model was obtained by averaging the bias scores for each attribute. The final bias scores showed that GPT-4o-mini was the most biased model, followed by ViT in the second place. Additionally, this table highlights that residual-based models, including ResNet and XceptionNet, showed the minimum bias. The final bias score for the POSTER and CEPrompt is not considerable, where they ranked as the fourth and third least bias models. Ultimately, our experiments rank the models from the most biased to the least biased as follows: GPT-4o-mini, ViT, CLIP, MobileNet, POSTER, CEPrompt, XceptionNet, and ResNet. Notably, this research focused solely on the FER task, and the observed bias in the models could be evaluated on different tasks, potentially yielding varying bias scores.
4.2.4 Conclusion
The analysis of the eight FER models demonstrates that all exhibit varying degrees of bias across Age, Gender, Race, and expression attributes. While specialized models such as POSTER and CEPrompt achieve higher accuracy, they do not eliminate bias. Overall, GPT-4o-mini and ViT were the most biased models, whereas ResNet and XceptionNet showed the least bias. These results highlight that high accuracy does not necessarily correspond to fairness, emphasizing the need to evaluate both performance and bias when assessing FER models.
5 Discussion and Future Works
Our analysis highlights that both the datasets and models used for facial expression recognition (FER) exhibit significant biases that could impact the fairness and generalizability of FER systems. Across the four in-the-wild datasets—AffectNet, Fer2013, RAF-DB, and ExpW—demographic imbalances are apparent. Race consistently emerges as the most biased attribute, while Age and Gender exhibit moderate but persistent disparities. Similarly, expression distributions are skewed, with Neutral and Happy overrepresented and Fear and Disgust underrepresented. These imbalances suggest that models trained on these datasets may inherit and even amplify existing biases.
The model evaluation further emphasizes the distinction between accuracy and fairness. While specialized FER models such as POSTER and CEPrompt achieve high overall accuracy, they do not effectively mitigate bias, particularly for underrepresented expressions and demographic groups. GPT-4o-mini and ViT, despite strong performance in some tasks, were identified as the most biased models, whereas residual-based architectures like ResNet and XceptionNet demonstrated relatively lower bias. Interestingly, some models partially compensated for imbalances in the data, reducing the impact of skewed distributions, but no model completely addressed the fairness challenges inherent to FER tasks.
These findings underline two important points. First, high accuracy alone is not sufficient to ensure equitable FER performance across all demographic groups and expressions. Second, addressing bias requires attention at both the dataset and model levels: collecting more balanced datasets, particularly for underrepresented expressions and demographic groups, is essential, as is developing models that explicitly account for these biases during training. Overall, our results call for a holistic approach that evaluates FER systems not only on predictive performance but also on fairness metrics, ensuring more reliable and equitable deployment in real-world applications.
The contributions of this research are pivotal for advancing fairness in FER tasks. Below, we outline several open research directions in FER that can build on this study to address bias and promote fair decision-making:
  • Investigating non-demographic sources of bias, such as illumination, background noise, head pose, gestures, eye gaze, and hair color, within datasets and their impact on fairness of models.
  • Reporting the bias score of trained models for FER tasks alongside their accuracy, which enables a comprehensive evaluation, ensuring both the effectiveness and fairness of the models.
  • Comparing the biases present in lab-controlled datasets versus in-the-wild datasets, and analyzing their influence on model training and fairness outcomes.
  • Exploring face generator models as potential sources of bias in future datasets and developing fair data generation methods to address these issues.
  • Developing more robust FER models by incorporating fairness constraints into the loss function during training. Future work could examine methods for implementing such constraints and their effects on model accuracy.
  • Analyzing the evolution of bias throughout the different stages of model development, including pretraining, training, fine-tuning, and domain adaptation.
  • Exploring the role of self-supervised learning methods in reducing bias within datasets and evaluating how they can contribute to fair model development.
  • Evaluating video-based datasets and models, where temporal dynamics between sequences are critical, to mitigate bias in video-based FER tasks.
  • Investigating bias in multi-modal models that leverage data from diverse sources such as images, text, video, audio, and physiological signals.
This research highlighted the significance of maintaining a balanced data distribution across demographic groups and ensuring the fairness of deep models. Establishing a foundation for enhancing the equity of FER models is essential. These improvements have the potential to benefit critical applications, including human-computer interaction (HCI), mental health monitoring, and surveillance. Future studies will focus on proposing methods to address bias in FER tasks while maintaining high accuracy.